# Effect of kinetic energy operator on double heterostructures spectra

### 24 Apr 2026 v1

@@ Abstract
Keywords— Double heterostructure, effective position-depend mass, von Ross’s operator, boundary conditions, kinetic energy operator, energy spectra, reflection and transmission coefficients.
@@ 1 Introduction
When considering heterostructures in condensed matter physics, the low-energy excitations of charge carriers are described by an effective Schrödinger equation, where the KEO is an ambiguous quantity. This ambiguity is due to the fact that the mass operator $m=M_{0}m(z)$, no longer commutes with the momentum operator $\hat{p}=-i\hbar d/dz$, leaving the called von Ross KEO [34]
@@
which has the ambiguity defined by conditions $\alpha+\gamma=-\frac{2}{3}$, $\alpha\gamma=0$ and therefore is a particular case of von Ross KEO with $\alpha\neq\gamma$. For this KEO, $T_{L}$, the boundary conditions for the wave function $\psi(z)$ of charge carriers at points $z_{k}$ within the heterostructure, are [23]
@@ 2 Schrödinger equation for position-dependent effective masses
Then, for appropriate $V(z)$ and $m(z)$, the associated Eq. (12) could be identified with a simpler or already known one. When this is the case, $E$ could be obtained directly and $\psi$ from $\phi$ and $m$ according to Eq. (11), where, in the cases of smooth profiles, the ‘generals’ boundary conditions must be taken into account as we will see in Section 3.1. This idea is one of the techniques already applied in Ref. [33], for two DHs that we again report in Secs. 4.1 and 4.2. The same idea has also been applied in Ref. [6] in two GHs to discard KEOs. In this work, each of these two GHs, is trapped in a DH (Eq. (2)), as described in Secs. 4.4, and 4.5. Given that from our DH models we can return to the GH model, we critically discuss the procedure followed in Ref [6] to discard KEOs.
@@ 3.1 First strategy: ‘analytic’ method
This is a transcendental equation that has been obtained through the analytical process described and allows us to calculate the energies of the bound states, as seen later in the DHs of Sections 4.1, 4.2, 4.4, and 4.5. The energy spectra of these analytical bound states are compared with the reflection coefficient poles which can be calculated using the multi-step methodology of Section 3.2.
@@ 4.4 Exponential potential and mass profiles
  • The term $g/\rho^{2}$ is a repulsive (attractive) potential barrier with a positive (negative) singularity at the origin only when its coefficient $g\geq 0$ ($g<0$). Nevertheless, this term retains the isochronous property that characterizes the harmonic oscillator when $g\geq-1/2$. This restriction includes potentials with $1/2\leq g<0$, where the intensity of the centripetal coupling is not strong enough to overcome the harmonic potential barrier. The energy spectrum of this isotonic potential is [4, 35, Cariñena_2008, 18, 30].
@@ 5 Conclusions
We have used the ordering of the kinetic energy operator that has been proposed in reference [23] and have also shown that von Ross’s KEO includes it. We have not found double heterostructures with ambiguities that prevent obtaining well-defined energy spectra, contrary to what has been reported in other references. For example, Ref. [31] establishes $\alpha=\gamma=0$ as the universally consistent ordering for KEO in effective mass theory, ensuring physically meaningful spectra for particle jumping in mass within a finite one-dimensional box. According to our results, we consider that the conclusions made in this reference are no longer valid for the type double heterostructures that we have adopted in this work. At the same time, we have critically discussed some of the arguments that were used for the conclusions made in the Ref. [6].
@@
RVT acknowledges Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI - México) support under the grant FORDECYT-PRONACES/61533/2020. JGR acknowledges IPN project SIP 20260578 and SECIHTI. JA acknowledges IPN project SIP20260576 and SECIHTI.

### 21 May 2026 v2

@@ Abstract
Keywords— Double heterostructure, effective position-depend mass, von Roos’s operator, boundary conditions, kinetic energy operator, energy spectra, reflection and transmission coefficients.
@@ 1 Introduction
When considering heterostructures in condensed matter physics, the low-energy excitations of charge carriers are described by an effective Schrödinger equation, where the KEO is an ambiguous quantity. This ambiguity is due to the fact that the mass operator $m=M_{0}m(z)$, no longer commutes with the momentum operator $\hat{p}=-i\hbar d/dz$, leaving the called von Roos KEO [34]
@@
which has the ambiguity defined by conditions $\alpha+\gamma=-\frac{2}{3}$, $\alpha\gamma=0$ and therefore is a particular case of von Roos KEO with $\alpha\neq\gamma$. For this KEO, $T_{L}$, the boundary conditions for the wave function $\psi(z)$ of charge carriers at points $z_{k}$ within the heterostructure, are [23]
@@ 2 Schrödinger equation for position-dependent effective masses
Then, for appropriate $V(z)$ and $m(z)$, the associated Eq. (12) could be identified with a simpler or already known one. When this is the case, $E$ could be obtained directly and $\psi$ from $\phi$ and $m$ according to Eq. (11), where, in the cases of smooth profiles, the ‘generals’ boundary conditions must be taken into account as we will see in Section 3.1. This idea is one of the techniques already applied in Ref. [33], for two DHs that we again report in Secs. 4.1 and 4.2. The same idea has also been applied in Ref. [6] in two GHs to discard KEOs. In this work, each of these two GHs, is trapped in a DH (Eq. (2)), as described in Secs. 48, and 4.5. Given that from our DH models we can return to the GH model, we critically discuss the procedure followed in Ref [6] to discard KEOs.
@@ 3.1 First strategy: ‘analytic’ method
This is a transcendental equation that has been obtained through the analytical process described and allows us to calculate the energies of the bound states, as seen later in the DHs of Sections 4.1, 4.2, 48, and 4.5. The energy spectra of these analytical bound states are compared with the reflection coefficient poles which can be calculated using the multi-step methodology of Section 3.2.
@@ 4.4 Exponential potential and mass profiles
  • The term $g/\rho^{2}$ is a repulsive (attractive) potential barrier with a positive (negative) singularity at the origin only when its coefficient $g>0$ ($g<0$). Nevertheless, this term retains the isochronous property that characterizes the harmonic oscillator when $g\geq-1/2$. This restriction includes potentials with $1/2\leq g<0$, where the intensity of the centripetal coupling is not strong enough to overcome the harmonic potential barrier. The energy spectrum of this isotonic potential is [4, 35, Cariñena_2008, 18, 30].
@@ 5 Conclusions
We have used the ordering of the kinetic energy operator that has been proposed in reference [23] and have also shown that von Roos’s KEO includes it. We have not found double heterostructures with ambiguities that prevent obtaining well-defined energy spectra, contrary to what has been reported in other references. For example, Ref. [31] establishes $\alpha=\gamma=0$ as the universally consistent ordering for KEO in effective mass theory, ensuring physically meaningful spectra for particle jumping in mass within a finite one-dimensional box. According to our results, we consider that the conclusions made in this reference are no longer valid for the type double heterostructures that we have adopted in this work. At the same time, we have critically discussed some of the arguments that were used for the conclusions made in the Ref. [6].
@@
RVT and ECO acknowledges Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI - México) support under the grant FORDECYT-PRONACES/61533/2020. JGR acknowledges IPN project SIP 20260578 and SECIHTI. JA acknowledges IPN project SIP20260576 and SECIHTI.